OpenStudy (anonymous):

I need some help interpreting an example problem using the \(\epsilon-\delta\) limit definition.

OpenStudy (anonymous):

Prove that \[\lim_{x \rightarrow 2}x^{2} = 4\] I know that you end up with \(|x – 2||x + 2| < \epsilon\) and that \(0 < |x – 2| < \delta\). Now, they say that for all x in the interval (1, 3), you know that |x + 2| <5. Then, letting \(\delta\) be the minimum of \(\frac{\epsilon}{5}\) and 1, it follows that, whenever 0 < |x – 2| < \(\delta\), you have \[|x^2 – 4| = |x – 2||x + 2| < (\frac{\epsilon}{5})(5) = \epsilon\]Now, I get everything up to there they say that x in the interval (1, 3) is true to |x + 2| < 5. I don't get anything afterwards. Can someone explain to me?

OpenStudy (anonymous):

Actually, I don't get the interval part either.

OpenStudy (anonymous):

can we start all over again?

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Ok, now you can start looking to the difinition of the limit , vwhich's \[\lim_{x \rightarrow x0} f(x)=L\]

OpenStudy (anonymous):

This is an example problem, so they just gave me the prove that \[\lim_{x \rightarrow 2}x^2 = 4\] using the \(\epsilon-\delta\) limit definition.

OpenStudy (anonymous):

now you probably know the two inequalities that we should form to find delta in temrs of epsilon, which are |x-2|<delta , and |\[x^{2}\] -4 | <epsilon

OpenStudy (anonymous):

Yes, that is what I am up to in my thinking process. I just can't figure out how to show that the \(\delta\) is possible.

OpenStudy (anonymous):

you just need to give at least one possible \[\delta\] , usually we find the maximum delta

OpenStudy (anonymous):

if one \[\delta\] is possible then the limit exists

OpenStudy (anonymous):

can I go on ?

OpenStudy (anonymous):

Yes. That's what I meant with my statement above if I was unclear.

OpenStudy (anonymous):

sorry it's just my English lol,, anyways, now we try to get rid of the absolute value in both inequalities

OpenStudy (anonymous):

we get -epsilon< \[x^{2}\] < epsilon ,, and ,, -delta< x-2< delta

OpenStudy (anonymous):

sorry it's epsilon<x^2-4< epsilon

OpenStudy (anonymous):

I feel im doing bad so far :(

OpenStudy (anonymous):

Ok, so you just split the absolute value into double inequalities right? No, you're doing well. :) Originally, I was trying to get where they got something from in the solution, but if there is a way of figuring it out this way, then it is great!

OpenStudy (anonymous):

aha you are right , so we now try to figure that out dude, just solve for x in both inequalities... we have: 4-epsilon <x^2 < 4+epsilon ,then, \[\sqrt{4-\epsilon}< \left| x \right| < \sqrt{4+\epsilon} \]

OpenStudy (anonymous):

now since x goes to a positive number which 's 2 , then we get \[\sqrt{4-\epsilon}< x < \sqrt{\epsilon+4}\]

OpenStudy (anonymous):

you can from this point consider two cases which are : 1) epsilon > 4 and i this case we can't use the current inequality we just come back to the one before taking root's . 2) epsilon <4 and in this case we use it ... am I clear so far

OpenStudy (anonymous):

I get up to where you found the roots. I still don't quite get why you needed to find those roots in the first place.

OpenStudy (anonymous):

to find two intervals where x belongs to,I mean one in terms of delta and the other in terms of epsilon so that we can make the endpoints equal to each other,, and choose the smallest delta between the two deltas I got after making the two endpoints equal to each other

OpenStudy (anonymous):

so the smallest delta I take between the two ,, we call it the maximum possible delta

OpenStudy (anonymous):

Ok, I think I understand now.

OpenStudy (anonymous):

OK, that's great,, now considering the (second case) , the inequalities I got are \[\sqrt{4-\epsilon} < x <\sqrt{4+\epsilon}\] and \[2-\delta < x < 2+\delta \]

OpenStudy (anonymous):

so take \[2-\delta = \sqrt{4-\epsilon} \] and \[2+\delta =\sqrt{4+\epsilon} \]

OpenStudy (anonymous):

I think now you can solve for delta :)

OpenStudy (anonymous):

So, then delta would be solved in terms of epsilon?

OpenStudy (anonymous):

yes, and you can say that delta = minimum of {2-sqrt(4-epsilon) , sqrt(4+epsilon)-2}

OpenStudy (anonymous):

remember that if only one delta exists , then the limit exists. and in our case there are infinite number of deltas

OpenStudy (anonymous):

Ok. Thank you. I'm a beginning calc student in self-study, so it's a bit hard to learn the material.

OpenStudy (anonymous):

dude don't forget about the first case lol

OpenStudy (anonymous):

self study is really helpful ,, keep up do not let anything let you down ,, I just finished calculus in the summer sission

OpenStudy (anonymous):

If I asked you to solve the first case on your own, will that be possibility ?

OpenStudy (anonymous):

Possibly. Just takes a while when they stump me like this.

OpenStudy (anonymous):

But I feel like I've taken up enough time lol. I'm sure once I take a break and come back to the problem, I can figure it out.

OpenStudy (anonymous):

okay , I will just give you a hint , before taking roots,, 4-epsilon is a negative number so the inequality will be automatically held ..

OpenStudy (anonymous):

I hope I was helpful for you, even though I believe I wasted alot of time while typing D: anyways If you couldn't do it just send me a msg or whatever u like thank u for your pationce

OpenStudy (anonymous):

Thank you for the help. :)